### v1

# Anti-aligning Self-propelled Model of Two Species: Emergence of Self-organized Heterogeneous Aligned and Clustered Order

Abstract
Self-propelled particles with anti-aligning interactions generally do not form a polar order. However, in this Letter, we show that when multiple types of such particles coexist and interact through aligning interactions between different species, a global polar order can emerge through the formation of elongated clusters with alternating domains of each species. By developing a mean-field theory, we reveal the conditions for cluster formation and characterize the resulting patterns. Our findings highlight the critical role of inter-species interactions in the emergence of complex ordered states.
Collective motion is widely observed in various biological systems, such as birds [1], fish [2], insects [3], bacteria [4], and cytoskeletal filaments [5, 6], and manifest self-organization in living systems. Such systems show a global polar order that spontaneously emerges through local interactions without external control. It is currently a widely studied topic in non-equilibrium statistical physics [7, 8, 9, 10]. The Vicsek model is the most pioneering and representative model [11], in which self-propelled particles align with their neighbors, leading to the formation of a polar order and collective movement in a cluster. Besides the Vicsek model, various other models have been proposed to study collective motion, most of which incorporate aligning interactions [12].
In contrast, even though the anti-aligning interaction has been recognized as a source of rich physical phenomena in systems other than those of active particles, e.g., the spin and neural network systems [13, 14], the behavior of active particles with anti-aligning interaction has not been sufficiently studied. One possible reason is the traditional belief that such interactions lead to a disordered state, that is, the particles move randomly and do not form a coherent structure. Indeed, recent studies have demonstrated that the anti-aligning Vicsek model (AAVM) does not show a global polar order across the entire system, although repetitive formation and destruction of hexagonal patterns of small particle clusters [15, 16] are observed. The extension of the AAVM to a two-species model with anti-aligning interactions between the species results in a global polar order within each species, but not across species [17]. These results challenge the traditional belief that anti-aligning interactions cannot lead to organized structures and raise the following question: How can self-propelled particles achieve a global polar order across an entire system despite the presence of anti-aligning interactions?
In this Letter, we introduce multiple species of anti-aligning self-propelled particles and demonstrate that a global polar order emerges through aligning interactions between different types of particles. In the steady state, the different particles form an elongated cluster with a stripe pattern. Within the cluster, the particles are aligned in the same direction along the long axis, causing the cluster to move ballistically along that axis. We term this polar and clustered state as the Self-organized Heterogeneous Aligned and Clustered Order (SHACO). For the emergence of the SHACO state, the aligning interaction between the different types of particles does not need to be larger than the anti-aligning interaction between the same type of particles. The stability of the SHACO state can be explained using the mean-field analysis, which can also effectively describe the characteristics of the stripe pattern and transition between the disordered and SHACO states. In previous studies of pattern formation in multi-species systems, non-reciprocal interactions between different particles were mainly investigated [18, 19, 20, 21, 22], but we found that even if the interaction is reciprocal, multi-species of self-propelled particles will show a characteristic pattern.
Model.— We consider self-propelled particles comprising two distinct populations, named type A and type B, in two dimensions. The particles either tend to align or anti-align with each other depending on their type. The number of self-propelled particles of type A and B are $N_{A}$ and $N_{B}$, respectively. The position of the $i$-th particle at time $t$ is represented by the vector $\bm{r}_{i}(t)=(x_{i}(t),y_{i}(t))$, and the particle moves in the direction $\theta_{i}(t)$ with a constant velocity subject to noise. The dynamics of each particle are the same as those in the time-continuous Vicsek model.
$\dot{\bm{r}_{i}}$ $=v_{0}\bm{s}_{i}+\sqrt{\sigma}\bm{\xi}_{i}\ ,$ (1)
$\tau\dot{\theta}_{i}$ $=-\langle m_{ij}\sin{(\theta_{i}-\theta_{j})}\rangle_{R}\ .$ (2)
Here ${\bm{s}}_{i}\equiv(\cos\theta_{i},\sin\theta_{i})$ is a unit vector representing the direction of self-propulsion, $\bm{\xi}_{i}$ represents a Gaussian white noise of zero mean and unit variance, and $\sigma$ is the dispersion of the noise. $v_{0}$ is the self-propelled velocity and is normalized to unity by selecting an appropriate time unit. The angle brackets $\langle\cdots\rangle_{R}$ in Eq. (2) denote the average over the neighboring particles $j$ that are within the distance of $R$ from the particle $i$. The angle $\theta_{i}$ tends to align or anti-align with the neighboring particles $j$ depending on the sign of $m_{ij}$: The $i$-th particle aligns (anti-aligns) with the $j$-th particle for a positive (negative) $m_{ij}$. $m_{ij}$ is set to $m_{AA}$ ($m_{BB}$) if both particles are of type A (type B) and to $m_{AB}$ or $m_{BA}$ if the particles in the interacting pair are of different types. $m_{AB}$ is generally distinct from $m_{BA}$ ($m_{AB}\neq m_{BA}$), but in this study we restricted our analysis to the symmetric case of $m_{AA}=m_{BB}$ and $m_{AB}=m_{BA}$, i.e., the system is reciprocal. $\tau$ is the ratio of the characteristic timescales of directional rotation to particle motion. Because we consider the anti-aligning interaction between the same type of particles, we can set $m_{AA}=m_{BB}=-1$ without loss of generality, and set $m_{AB}=m_{BA}=c$, where $c$ is in the range of $0\leq c\leq 1$. We term this model as the two-species AAVM. For comparison, we also conducted a numerical simulation of the one-species AAVM [15]; see the supplementary material [23].
We numerically simulated the above Eqs. (1) and (2) in a two-dimensional $L\times L$ square region under periodic boundary conditions. The parameters were set as $v_{0}=1.0,\tau=0.05$, $R=3.0$, $N_{A}=1,000$, $N_{B}=1,000$, $L=32$, and the noise strength $\sigma=0.04$, unless otherwise noted. The initial conditions were randomly chosen such that both the position ${\bm{r}}_{i}$ and angle $\theta_{i}$ were uniformly distributed.
Figure 1: Behavior of the two-species AAVM at $c=1.0$. (a) Snapshots at early, middle, and late stages (from top to bottom). Type A and B particles are represented by red and blue dots, respectively. The black dashed circle indicates the interaction region of radius $R$. The black arrow indicates the forward direction of the cluster. (b) The mean squared displacement (MSD) of a particle in the steady state. Green, blue, and dashed black lines indicate the MSDs of the two-species AAVM, single-species AAVM, and Brownian motion, respectively. (c, d) Mean particle density of a single particle within the interaction radius $R$. Color maps indicate the particle density of (c) the same-type and (d) different-type particles in the steady state. The coordinates were chosen such that the center particle moves in the positive direction of the abscissa. The density is normalized by the average density of the particles within the distance $R$ (the number of type A or B particles within the interacting region divided by the area $\pi R^{2}$). The regions with higher and lower densities than the averaged density are represented by yellow- and black-color scales, respectively.
Results.— Figure 1(a) shows the observed time series of the two-species AAVM with the parameter $c=1.0$ (from top to bottom). Starting from the random disordered state, the particles form several small clusters with a striped pattern. These clusters collide and merge with each other repeatedly, forming one large cluster in the steady state (Movie SMov. 1 in [23]). The cluster moves in one direction. Thus, the mean squared displacement (MSD) of the particles is proportional to $\Delta t^{2}$ indicating ballistic behavior (Fig. 1(b) green line). This contrasts with the long-term dynamics of the one-species AAVM that exhibits diffusive behavior (Fig. 1(b) blue line and Supplementary Fig. 1 in [23]).
In the cluster, each stripe consists of particles of the same type, indicating that particles of the same type aggregate and align despite their anti-aligning interactions. These stripes, composed of either type A or B particles, alternate in sequence parallel to the direction of cluster propulsion. Notably, the stripe spacing $\ell$ remains nearly constant throughout the cluster and is proportional to $R$; $\ell/R$ is $0.615\pm 0.015$, calculated through statistical analysis of $R$ incremented from $1.5$ to $3.5$ in steps of $0.5$, where the ratio $\ell/R$ is almost independent of the interaction radius. In addition, the width of the cluster in the steady state is approximately $2R$. When the total number of particles is larger, instead of the width increasing, the particle density within the cluster becomes greater; alternatively, the length of the cluster is extended by increasing the number of stripes while maintaining a stripe spacing.
Reflecting the equispaced stripe patten, the average particle density within the interaction radius of a reference particle also exhibits a stripe pattern (Fig. 1(c) and (d)). Particles of the same type or different types are highly concentrated within the striped domains, which are separated by a distance $\ell$ and do not overlap. This stripe pattern in the average density distribution remains unchanged and scales only when $R$ is changed.
These observations indicate that the inter-species interaction between the different types of anti-aligned self-propelled particles leads to a global polar order with a striped cluster. Because such a polar state with a striped cluster has not been observed either in the one-species AAVM (see [15] and Supplementary Fig. 1 in [23]) or in the original Vicsek model and its variants according to the aligning interaction (see [11]), we propose the term SHACO [24] for this state.
Figure 2: Polar order parameter dependence on the parameter sets of (a) $(c,R)$ and (b) $(c,\tau)$. Color map indicates the polar order parameter $\phi$. The black dashed line in (a) indicates the analytically calculated transition point $c=0.69$ (see the main text). We used $L=64$ as the system size for calculating the order parameter. The other parameters were fixed to the values described in the main text. (c) Snapshots of the simulations using the parameters indicated by the symbols in (b): $(c,\tau)=(0.9,0.5)$ (red triangle), $(0.9,2.81\times 10^{-2})$ (green square), and $(0.9,1.58\times 10^{-3})$ (blue circle). (d) Number fluctuation measured using the time-averaged particle number $\langle N\rangle$ and its standard deviation $\Delta N=\sqrt{\left(N-\langle N\rangle\right)^{2}}$ within a given observation region. $\Delta N$ scales asymptotically as $\langle N\rangle^{\alpha}$. The same simulation data as in (b) and (c) were used.
In which parameter region, can this SHACO state be observed? We measured the polar-order parameter $\phi\equiv\left|(1/N)\sum_{j}e^{i\theta_{j}}\right|$ as an indicator of the SHACO phase in various ranges of the parameter sets of $c$, $R$, and $\tau$ (Fig. 2). In the short interaction radius limit $R\rightarrow 0$, the aligning interaction between different types of particles $c$ must be as strong as the anti-aligning interaction between the same type of particles ($m_{AA}=m_{BB}=-1$) for the emergence of the SHACO state. The required strength of $c$ decreases as $R$ increases, finally reaching the plateau at $c\simeq 0.7$ as indicated by the black dashed line in Fig. 2(a). This suggests that even if the magnitude of interactions between different types of particles does not exceed that between the same type of particles, the heterogeneous particles can facilitate the emergence of the polar order.
Even when $c$ is sufficiently large, $\tau$ must be in the specified range for the emergence of the SHACO state (Fig. 2(b)). In regions where $\tau$ is either too large or small, the polar order disappears in the steady state. Although these two states cannot be distinguished in the polar order parameter, they exhibit different behaviors in terms of the number fluctuation (Fig. 2(c) an (d); Movie SMov. 2 in [23]). In the region with a small $\tau$, the particle density largely fluctuates locally and the exponent of the number fluctuation, $\alpha$, is approximately $0.8$, which is similar to the value observed in the Vicsek model and is known as the giant number fluctuation (GNF) [25, 26, 27]. In the region with a large $\tau$, the exponent is approximately $0.5$, as in the normal fluctuation. This suggests that the phases with small and large $\tau$ are distinct, and in the small-$\tau$ phase, the characteristics of the Vicsek model are recovered even in the presence of anti-aligning interactions. The range of $\tau$ for stabilizing the SHACO state increases as $c$ becomes larger. Fig. 2(b) suggests that the upper and lower critical values of $\tau$ are approximately proportional to $\exp(\beta_{\pm}c)$, where $\beta_{\pm}$ denotes the positive or negative constants for the upper and lower transition boundaries.
Figure 3: Mean-field analysis of the interaction energy. (a) Distribution of particles assumed by idealizing the numerically obtained particle density (Fig. 1 (c) and (d)). Yellow and gray areas represent the regions where the same-type and different-type particles exist in high concentrations, respectively. (b) Interaction energy normalized by the area within the interaction radius, $E(\ell)/R^{2}$, calculated through the mean-field analysis. We calculated $E(\ell)/R^{2}$ by setting $c$ to $1.0$ (blue) and $0.69$ (green). The red dashed line corresponds to $\ell/R=2/\sqrt{10}$ at which $E(\ell)$ is the minimum irrespective of $c$.
Why is the stripe pattern of the SHACO state stable, and why does the transition to the SHACO state occur even when the magnitude of the interaction between particles of different types does not exceed that between particles of the same type? In the stable striped cluster, each stripe consists of almost only one type of particle, and the average particle distribution forms a regular striped pattern in the interaction radius of $R$, as shown in Fig. 1(c) and (d). This observation allows us to develop a mean-field theory to address the condition for the emergence of the polar order. We assume that the idealized particle distribution within the interaction radius of $R$ is the stripe of a different species with a constant spacing $\ell$, as illustrated in Fig. 3(a). In addition, we employ the mean-field assumption in which the particle concentration is uniform in each stripe and all particles are exposed to the same particle field. For a given value of $\ell$, the interaction region is divided into $12n_{s}+$ 1 striped domains, where $12n_{s}+$ 1 is the smallest odd integer greater than or equal to $2R/\ell$. All particles are assumed to be aligned in the positive direction of the $x$-axis. In this mean-field distribution, the direction $\theta$ of the particle at the center follows $\tau\dot{\theta}=E(\ell)\sin\theta$, with the interaction energy $E(\ell)$ given by
[…]
where $\sigma_{\ell}(x)=-1$ in the range of $(2n-1/2)\ell<x<(2n+1/2)\ell$, where $n$ is an integer, and $\sigma_{\ell}(x)=c$ otherwise.
The functional form of $E(\ell)$ is shown in Fig. 3(b). $E(\ell)$ takes its minimum at $\ell=2R/\sqrt{10}\approx 0.632R$, which coincides with the observed stripe spacing in the clusters (see the supplementary material [23]). This result demonstrates the smallest value of $E(\ell)$ corresponding to the most stable striped pattern was realized in the simulations. The spacing is proportional to $R$ and is independent of the other parameters. In addition, the minimum value of $E(\ell)$ must be negative for the polar order to be stable. Because the minimum value is calculated as $E(2R/\sqrt{10})=2R^{2}\left(\cos^{-1}\!\left(4/5\right)-c\sin^{-1}\!\left(4/5\right)\right)$, the transition point between the SHACO state and disordered state is evaluated as $c\approx 0.69$, which is in excellent agreement with the minimum value of the numerically obtained $c$ for the polar order to emerge (Fig. 2(a), black dashed line). Note that, when $R$ is small, the fluctuations in the number of particles within the interaction radius are too large to ignore. Then, the finite-size effects cause the transition point $c$ to shift to higher values, as the reduced correlation length makes it more difficult for the particles to align.
Discussion.— In this Letter, we report a novel global polar-ordered state, termed the SHACO state, incorporating multiple species of anti-aligning self-propelled particles. The emergence of the SHACO state critically depends on the interaction strength between different types of particles, denoted by $c$, and timescale of alignment, $\tau$. Remarkably, the SHACO state can appear even when the aligning interaction between different types of particles is weaker than the anti-aligning interaction among particles of the same type, achieved through the formation of striped clusters. The mean-field analysis explains the structure of the cluster and effectively captures the parameter dependence on $c$, while the dependency on $\tau$ remains an open question. Upon examining the polar order parameter $\phi$, it appears that the SHACO state undergoes a reentrant transition with variation in $\tau$. However, two phases with too large or small $\tau$ can be distinguished by the emergence of a GNF (Fig. 2(d)). Understanding why the GNF, similar to that observed in the original Vicsek model, is observed even in the presence of anti-aligning interactions, and how the transition between the GNF and clustered phases—absent in the Vicsek model—occurs, are potential topics for future research.
Recent studies have demonstrated that multi-species systems of self-propelled particles give rise to complex patterns [20, 21, 28, 29], including moving stripe patterns under reciprocal [30] and non-reciprocal [31, 32] interactions. However, these stripe patterns do not form clusters, and the stripes are perpendicular to the direction of motion, consistent with the fact that self-propelled particles with aligning interactions typically form moving bands that extend perpendicular to their direction of motion known as Vicsek waves [27, 33, 34, 35]. In contrast, the moving cluster observed in this study elongates along the direction of motion while maintaining a constant width. This unique behavior is attributed to the conflicting interactions between particles of the same type and those of different types. Although a comprehensive analysis, including that based on theoretical approaches such as the application of Boltzmann-Ginzburg-Landau theory [36], remains an open challenge, these conflicting interactions provide fresh insights into the dynamics of species self-propelled particles and will generate various new and complex patterns.
Such conflicting interactions are expected to be observed at various levels in living systems; for example, regarding animal behaviors, individuals of the same sex repel each other, whereas individuals of the opposite sex attract each other, for enhancing the chances of mating. Although it is difficult to identify the SHACO state in nature, it is much easier and more interesting to artificially design the interaction and control the state. Recent studies succeeded in synthesizing active matter with repulsive interactions [37, 38, 20]. The introduction of different types of particles and the conflicting interactions between them are expected to lead to the emergence of the SHACO state. A binary mixture system of programmable robots that analyzes neighboring individuals and controls their interactions (cf. Ref. [39]) is another candidate.
In conclusion, the multiple species of self-propelled particles and conflicting interactions between them give rise to new phenomena of active matter. We hope that further studies, including the exploration of experimental realizations discussed above, will enrich our understanding and provide deeper insights into the underlying mechanisms of non-equilibrium systems.
Acknowledgements.
We are grateful to J. Yamagishi and K. Mitsumoto for their fruitful discussions. This study was supported by JSPS KAKENHI Grant Numbers 24H01931 (to S.I.), 21K15048 (to T.S.H.), and JST CREST Japan, Grant Number JPMJCR1923 (to S.I.).